OpenStudy (anonymous):

What is 2x -3y = 12 in slope intercept form?

OpenStudy (kobeni-chan):

Slope intercept form is y=mx+b Right now, this is in standard form, ax + by = c You need to get y on one side of the equal sign by itself. 2x - 3y = 12 Add 3y to each side, this will make it 2x = 12 + 3y Subtract 12 from each side. 2x - 12 = 3y To isolate y, divide both sides by 3. (2/3)x - 4 = y You can switch it around to y = (2/3)x - 4

OpenStudy (anonymous):

2x - 3y = 12 -2x -2x -3y = -2x +12 -3y/-3 -2/-3 12/-3 Y= 2/3x -4

OpenStudy (anonymous):

So the slope is 2/3 and the y intercept is -4?

OpenStudy (anonymous):

Thank you!